# DeepResonance: Enhancing Multimodal Music Understanding via Music-centric Multi-way Instruction Tuning

### v1

@@ 5.5 Ablation Study
Figs. 5 and 6 present the results of ablation studies evaluating the effectiveness of key components in the proposed methods, including Music4way-MI2T and Music4way-MV2T instruction tuning data (MWIT), multi-sampled ImageBind embeddings (MIE), and the pre-alignment Transformer (PT). For music + text $\xrightarrow{}$ text tasks (Fig. 6), we observe that MWIT, MIE, and PT each contribute positively to performance. However, when combined, PT does not consistently complement the other two components across all three benchmarks, with MWIT + MIE (DeepResonance-$\alpha$) yielding the consistent improvements. For multimodal music understanding tasks (music + image/video + text → text), we compare settings with and without MIE and PT in Fig. 5, as MWIT serves as the in-domain data for these tasks. Integrating MIE and PT enhances performance, with PT proving most effective when limited to a single Transformer layer. This highlights the effectiveness of MIE and PT while suggesting that increasing PT’s parameters may lead to overfitting on limited instruction tuning data. Moreover, PT with different layer settings negatively impacts performance on Music4way-Any2T (refer to Sec. 5.3).
@@ 6 Conclusion
    The input music training data is mostly limited to clips shorter than 30 seconds, with a significant portion (e.g., AudioSet) being 10 seconds. This may restrict the fine-tuned models’ performance on longer music sequences. Additionally, the dataset’s image frames are directly extracted from videos, assuming relevance within short clips (10s). For longer videos, selecting the most representative frames, such as cover images, should be explored in future work.
  • (2)
    The effectiveness of DeepResonance has been verified using Vicuna (based on LLaMA) as the backbone LLM, but further studies are needed to assess its improvements when applied to other LLM architectures.
  • (3)
    The extracted music features (e.g., downbeats, tempo, key, chords) rely on existing MIR algorithms, which may have precision limitations. Incorporating more signal-oriented features such as MFCCs or Chromagrams could be explored in future work.
@@ Appendix B Training and Evaluation Details
Following the training strategy of NExT-GPT and M2UGen, we train DeepResonance in two stages. In the first stage, we fine-tune only the parameters of the linear adaptors and the proposed pre-alignment Transformer. This stage focuses on captioning tasks for music, image, and video modalities, utilizing images from COCO Lin et al. (2014) and music and video clips from the constructed Music4way dataset. [^6] We use COCO instead of Music4way for the image captioning task in the first stage, as empirical results indicate that using the larger image-text dataset, COCO, yields better performance. In the second stage, we fine-tune the linear adaptors and the pre-alignment Transformer while simultaneously performing LoRA-based fine-tuning Hu et al. (2022) on Vicuna. This stage incorporates instruction tuning tasks using datasets including Alpaca Taori et al. (2023), MusicCaps Agostinelli et al. (2023), and MusicQA Liu et al. (2024) [^7] We use the “fine-tune” split of the MusicQA dataset and exclude the “train” split to avoid overlap with the test split of MusicCaps., along with our constructed Music4way, Music4way-MI2T, and Music4way-MV2T datasets. A summary of all datasets used for training and evaluation is provided in Table 1. As depicted in Fig. 2, instructions are fed directly into the text-LLM, bypassing the adaptors and the pre-alignment Transformer, as they do not require interaction with the input information.
@@
We fine-tune for $5$ and $2$ epochs in the first and second stages, respectively, utilizing a learning rate of $1\mathrm{e}{-4}$ and a batch size of $16$. Training is conducted on $8$ NVIDIA A100 GPUs (40GB each). For LoRA, the rank and alpha are both set to $32$, following NExT-GPT. We train the pre-alignment Transformer with various layer configurations (see Sec. 5.5) and find that a single Transformer layer achieves the best performance. Regarding the trainable parameters, the linear adaptors, LoRA, pre-alignment Transformer, and LLaMA embedding layers contain $4$M, $33$M, $157$M, and $262$M parameters, respectively, comprising 5.6% of the total model parameters. Regarding the training budget, stage 1 took $25.2$ hours, while stage 2 took $20.3$ hours for DeepResonance-$\alpha$. For DeepResonance-$\beta$, the training times were $27.0$ hours for stage 1 and $22.7$ hours for stage 2.
@@
For evaluation, we report the mean results from three inference runs and include BLEU Papineni et al. (2002), ROUGE-L Lin (2004), and BERTScore Zhang et al. (2020), following the setup of M2UGen Hussain et al. (2023). We report BLEU-1, BLEU, ROUGE-L precision, ROUGE-L recall, ROUGE-L F1, BERTScore precision, BERTScore recall, and BERTScore F1 details in Appx. EFG, and H.
Appendix C Construction Details and Data Examples of Music4way-MI2T and Music4way-MV2T
@@
Appendix D Construction Details and Data Examples of Music4way-Any2T
@@
Appendix E Detailed Results on Music Understanding Tasks
Tables 910, and 11 report the detailed performance on all metrics of all models introduced in Sec.5.1 on music understanding tasks (music + text $\xrightarrow{}$ text) using MusicQALiu et al. (2024), MusicCaps Agostinelli et al. (2023), and our constructed Music4way-MusicCaps.
Appendix F Detailed Results on Multimodal Music Understanding Tasks
Tables 1213, and 14 list the result details on all metrics on our newly proposed multimodal music understanding tasks (music + image/video + text $\xrightarrow{}$ text), including Music4way-MI2T, Music4way-MV2T, and Music4way-Any2T.
Appendix G Detailed Results of Zero-shot Evaluation
The detailed results of the zero-shot evaluation are presented in Tables 15161718, and 19.
Appendix H Detailed Results of Ablation Study
The detailed results of the ablation study are presented in Tables 2021222324 and 25.
@@ Appendix H Detailed Results of Ablation Study
Table 9: Results on MusicQA. The top two performances are highlighted in bold. “” indicates that the test set’s train split is included in the corresponding model’s training data.
@@
Table 10: Results on MusicCaps. The top two performances are highlighted in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set. “” indicates that the test set’s train split is included in the corresponding model’s training data.
@@
Table 11: Results on Music4way-MusicCaps. The top two performances are highlighted in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set. “” indicates that the test set’s train split is included in the corresponding model’s training data.
@@
DeepResonance-$\alpha$ (ours) 44.6 9.2 27.8 33.8 30.1 89.1 89.6 89.4
@@
Table 12: Results on Music4way-MI2T. The top two performances are highlighted in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set.
@@
DeepResonance-$\alpha$ (ours) 44.5 8.9 28.2 32.7 29.6 89.0 89.4 89.2
@@
Table 13: Results on Music4way-MV2T. The top two performances are highlighted in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set.
@@
DeepResonance-$\alpha$ (ours) 54.7 14.2 35.9 34.6 35.1 91.2 90.6 90.9
@@
Table 14: Results on Music4way-Any2T. The top two performances are highlighted in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set.
@@
Table 20: Ablation study on MusicQA. The top two performances are highlighted in bold.
@@
Table 21: Ablation study on MusicCaps. The top two performances are highlighted in bold.
@@
Table 22: Ablation study on Music4way-MusicCaps. The top two performances are highlighted in bold.

### v2

@@ 5.5 Ablation Study
Figs. 5 and 6 present the results of ablation studies evaluating the effectiveness of key components in the proposed methods, including Music4way-MI2T and Music4way-MV2T instruction tuning data (MWIT), multi-sampled ImageBind embeddings (MIE), and the pre-LLM fusion Transformer (PT). For music + text $\xrightarrow{}$ text tasks (Fig. 6), we observe that MWIT, MIE, and PT each contribute positively to performance. However, when combined, PT does not consistently complement the other two components across all three benchmarks, with MWIT + MIE (DeepResonance-$\alpha$) yielding the consistent improvements. For multimodal music understanding tasks (music + image/video + text → text), we compare settings with and without MIE and PT in Fig. 5, as MWIT serves as the in-domain data for these tasks. Integrating MIE and PT enhances performance, with PT proving most effective when limited to a single Transformer layer. This highlights the effectiveness of MIE and PT while suggesting that increasing PT’s parameters may lead to overfitting on limited instruction tuning data. Moreover, PT negatively impacts performance on Music4way-Any2T (refer to Sec. 5.3).
Model MusicQA MusicCaps Music4way-MusicCaps
B-1 R-L BERTS B-1 R-L BERTS B-1 R-L BERTS
DR-$\alpha$ 35.1 40.8 91.6 26.0 21.6 87.3 40.9 48.4 93.3
w/o v.c. 35.0 40.4 91.4 25.5 21.2 87.2 40.7 48.1 93.2
w/o m.f. 34.9 40.3 91.5 25.9 21.3 87.3 40.5 48.0 93.2
DR-$\beta$ 35.6 41.1 91.6 25.8 21.6 87.3 39.9 47.8 93.2
w/o v.c. 35.2 40.7 91.5 25.8 21.2 87.2 39.5 46.7 93.0
w/o m.f. 35.2 40.2 91.4 25.8 21.1 87.3 40.0 48.0 93.2
Table 4: Ablation study on the impact of decoder-side visual and musical contents. See details in Appx. J. “v.c.”: visual captions; “m.f.”: low-level musical features.
We also conduct an ablation study on the multi-way unified captions of Music4way-MI2T and Music4way-MV2T. Referring to Fig.3, we remove image or video captions to isolate the contribution of visual contents, and similarly remove low-level musical features to assess the impact of detailed musical information. As shown in Table 4, removing either leads to a slight performance drop, suggesting that the unified target formats of both contents enhance music understanding capability. However, comparisons in Fig.6 indicate that using MWIT and applying our proposed MIE or PT methods have a greater impact than the specific target-side content of the captions.
@@ 6 Conclusion
    The input music training data is mostly limited to clips shorter than 30 seconds, with a significant portion (e.g., AudioSet) being 10 seconds. This may restrict the fine-tuned models’ performance on longer music sequences. Additionally, the dataset’s image frames are directly extracted from videos, assuming relevance within short clips (10s). For longer videos, selecting the most representative frames—such as cover images—should be explored in future work, once more licensed long-form music and video clips become available. (This is beyond the scope of this paper, as scene transitions within the 10-second AudioSet clips are rare; therefore, randomly selecting a frame is sufficient.)
@@ Appendix B Training and Evaluation Details
Following the training strategy of NExT-GPT and M2UGen, we train DeepResonance in two stages. In the first stage, we fine-tune only the parameters of the linear adaptors and the proposed pre-LLM fusion Transformer. This stage focuses on captioning tasks for music, image, and video modalities, utilizing images from COCO Lin et al. (2014) and music and video clips from the constructed Music4way dataset. [^5] We use COCO instead of Music4way for the image captioning task in the first stage, as empirical results indicate that using the larger image-text dataset, COCO, yields better performance. In the second stage, we fine-tune the linear adaptors and the pre-LLM fusion Transformer while simultaneously performing LoRA-based fine-tuning Hu et al. (2022) on Vicuna. This stage incorporates instruction tuning tasks using datasets including Alpaca Taori et al. (2023), MusicCaps Agostinelli et al. (2023), and MusicQA Liu et al. (2024) [^6] We use the “fine-tune” split of the MusicQA dataset and exclude the “train” split to avoid overlap with the test split of MusicCaps., along with our constructed Music4way, Music4way-MI2T, and Music4way-MV2T datasets. A summary of all datasets used for training and evaluation is provided in Table 1. As depicted in Fig. 2, instructions are fed directly into the text-LLM, bypassing the adaptors and the pre-LLM fusion Transformer, as they do not require interaction with the input information.
@@
We fine-tune for $5$ and $2$ epochs in the first and second stages, respectively, utilizing a learning rate of $1\mathrm{e}{-4}$ and a batch size of $16$. Training is conducted on $8$ NVIDIA A100 GPUs (40GB each). For LoRA, the rank and alpha are both set to $32$, following NExT-GPT. We train the pre-LLM fusion Transformer with various layer configurations (see Sec. 5.5) and find that a single Transformer layer achieves the best performance. Regarding the trainable parameters, the linear adaptors, LoRA, pre-LLM fusion Transformer, and LLaMA embedding layers contain $4$M, $33$M, $157$M, and $262$M parameters, respectively, comprising 5.6% of the total model parameters. Regarding the training budget, stage 1 took $25.2$ hours, while stage 2 took $20.3$ hours for DeepResonance-$\alpha$. For DeepResonance-$\beta$, the training times were $27.0$ hours for stage 1 and $22.7$ hours for stage 2.
@@
For evaluation, we report the mean results from three inference runs and include BLEU Papineni et al. (2002), ROUGE-L Lin (2004), and BERTScore Zhang et al. (2020), following the setup of M2UGen Hussain et al. (2023). We report BLEU-1, BLEU, ROUGE-L precision, ROUGE-L recall, ROUGE-L F1, BERTScore precision, BERTScore recall, and BERTScore F1 details in Appx. GHI, and J.
Appendix C Existing Baseline Models
Appendix D Sanity check during the Inference Phase
We perform a sanity check during the inference phase to assess the contribution of multimodal inputs to overall performance. Specifically, we retain only the textual inputs while removing all other modalities—music, image, and video—and evaluate the model on all six supervised music understanding tasks. The results, presented in Table 6, show a clear performance drop when multimodal inputs are excluded. This suggests that the DeepResonance models effectively leverage semantic information from non-text modalities. Moreover, the observed performance degradation highlights the utility and validity of our constructed Music4Way evaluation datasets for testing multimodal music LLMs.
Appendix E Construction Details and Data Examples of Music4way-MI2T and Music4way-MV2T
@@
Appendix F Construction Details and Data Examples of Music4way-Any2T
@@
Appendix G Detailed Results on Music Understanding Tasks
Tables 1112, and 13 report the detailed performance on all metrics of all models introduced in Sec.5.1 on music understanding tasks (music + text $\xrightarrow{}$ text) using MusicQALiu et al. (2024), MusicCaps Agostinelli et al. (2023), and our constructed Music4way-MusicCaps.
We also report the results of DeepResonance models based on LLaMA-3888We used LLaMA-3-8B-Instruct. Dubey et al. (2024) and Mistral [^9] We used Ministral-8B-Instruct-2410. Jiang et al. (2023) in Tables 11,12, and13. Our default DeepResonance models built on Vicuna LLMs achieve the best overall performance, likely due to their smaller embedding layers, which are easier to fine-tune and adapt to multimodal contexts. Nevertheless, the Mistral- and LLaMA-3-based DeepResonance models still outperform their respective NExT-GPT counterparts trained on the Music4way dataset.
Appendix H Detailed Results on Multimodal Music Understanding Tasks
Tables 1415, and 16 list the result details on all metrics on our newly proposed multimodal music understanding tasks (music + image/video + text $\xrightarrow{}$ text), including Music4way-MI2T, Music4way-MV2T, and Music4way-Any2T.
Appendix I Detailed Results of Zero-shot Evaluation
The detailed results of the zero-shot evaluation are presented in Tables 17181920, and 21.
Appendix J Detailed Results of Ablation Study
The detailed results of the ablation study are presented in Tables 2223242526 and 27.
@@ Appendix I Detailed Results of Zero-shot Evaluation
NExT-GPT-Mistral w/ Music4way 33.9 14.1 50.3 37.2 39.7 92.2 90.3 91.2
DeepResonance-$\alpha$-Mistral (ours) 35.0 14.7 50.0 38.5 40.4 92.3 90.6 91.4
DeepResonance-$\beta$-Mistral (ours) 34.6 14.6 50.9 38.0 40.5 92.4 90.5 91.4
NExT-GPT-LLaMA-3 w/ Music4way 32.3 13.1 50.3 36.0 38.8 92.3 90.2 91.2
DeepResonance-$\alpha$-LLaMA-3 (ours) 33.9 13.9 49.5 37.6 39.5 92.1 90.4 91.2
DeepResonance-$\beta$-LLaMA-3 (ours) 33.3 13.8 50.4 36.8 39.4 92.2 90.2 91.2
Table 11: Results on MusicQA. The top two performances in the first block are highlighted in bold. Scores where Mistral or LLaMA-3 based models surpass their respective baselines are shown in bold. “” indicates supervised evaluation settings.
@@
NExT-GPT-Mistral w/ Music4way 25.3 2.7 22.6 21.1 20.8 87.4 86.9 87.1
DeepResonance-$\alpha$-Mistral (ours) 25.8 2.8 23.1 21.6 21.3 87.5 87.0 87.2
DeepResonance-$\beta$-Mistral (ours) 25.6 2.6 22.5 21.4 21.0 87.4 87.0 87.2
NExT-GPT-LLaMA-3 w/ Music4way 25.4 2.7 22.8 21.2 21.0 87.4 86.8 87.1
DeepResonance-$\alpha$-LLaMA-3 (ours) 26.0 2.7 23.0 21.6 21.3 87.5 87.0 87.2
DeepResonance-$\beta$-LLaMA-3 (ours) 25.8 2.8 23.1 21.6 21.3 87.5 87.0 87.2
Table 12: Results on MusicCaps. The top two performances in the first block are highlighted in bold. Scores where Mistral or LLaMA-3 based models surpass their respective baselines are shown in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set. “” indicates supervised evaluation settings.
@@
NExT-GPT-Mistral w/ Music4way 37.9 16.9 56.9 44.5 46.2 93.8 92.2 93.0
DeepResonance-$\alpha$-Mistral (ours) 39.1 18.6 59.4 45.1 47.9 94.2 92.2 93.2
DeepResonance-$\beta$-Mistral (ours) 39.7 18.9 58.4 45.7 47.8 94.0 92.4 93.2
NExT-GPT-LLaMA-3 w/ Music4way 37.9 16.9 56.9 44.5 46.2 93.8 92.2 93.0
DeepResonance-$\alpha$-LLaMA-3 (ours) 39.5 19.2 60.3 45.4 48.4 94.4 92.3 93.3
DeepResonance-$\beta$-LLaMA-3 (ours) 39.7 19.2 60.3 45.2 48.4 94.4 92.3 93.3
Table 13: Results on Music4way-MusicCaps. The top two performances in the first block are highlighted in bold. Scores where Mistral or LLaMA-3 based models surpass their respective baselines are shown in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set. “” indicates supervised evaluation settings.
@@
DeepResonance-$\alpha$ (ours) 48.7 16.4 37.1 36.8 36.2 90.3 90.2 90.2
@@
NExT-GPT-Mistral w/ Music4way 8.6 0.8 45.6 12.6 18.7 89.6 85.2 87.3
DeepResonance-$\alpha$-Mistral (ours) 48.9 16.6 37.8 36.9 36.7 90.4 90.2 90.3
DeepResonance-$\beta$-Mistral (ours) 49.8 17.5 37.5 38.3 37.2 90.3 90.4 90.4
NExT-GPT-LLaMA-3 w/ Music4way 5.7 0.2 37.7 10.4 15.0 86.9 83.1 84.9
DeepResonance-$\alpha$-LLaMA-3 (ours) 48.9 16.5 37.5 36.8 36.4 90.3 90.2 90.3
DeepResonance-$\beta$-LLaMA-3 (ours) 49.4 17.3 37.3 38.2 37.0 90.3 90.4 90.3
Table 14: Results on Music4way-MI2T. The top two performances in the first block are highlighted in bold. Scores where Mistral or LLaMA-3 based models surpass their respective baselines are shown in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set.
@@
DeepResonance-$\alpha$ (ours) 48.9 16.6 37.4 37.0 36.5 90.3 90.2 90.2
@@
NExT-GPT-Mistral w/ Music4way 9.0 0.9 46.1 12.8 19.0 89.7 85.3 87.4
DeepResonance-$\alpha$-Mistral (ours) 48.7 16.4 37.5 36.7 36.4 90.4 90.2 90.3
DeepResonance-$\beta$-Mistral (ours) 49.7 17.5 37.5 38.3 37.2 90.3 90.4 90.4
NExT-GPT-LLaMA-3 w/ Music4way 4.9 0.2 38.5 10.5 14.9 86.8 83.1 84.9
DeepResonance-$\alpha$-LLaMA-3 (ours) 49.0 16.7 37.6 36.8 36.5 90.4 90.2 90.3
DeepResonance-$\beta$-LLaMA-3 (ours) 49.5 17.5 37.2 38.2 36.9 90.2 90.4 90.3
Table 15: Results on Music4way-MV2T. The top two performances in the first block are highlighted in bold. Scores where Mistral or LLaMA-3 based models surpass their respective baselines are shown in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set.
@@
DeepResonance-$\alpha$ (ours) 37.2 6.0 36.0 25.7 29.6 90.8 88.3 89.5
@@
NExT-GPT-Mistral w/ Music4way 17.4 1.8 41.8 17.8 24.0 90.0 86.2 88.0
DeepResonance-$\alpha$-Mistral (ours) 37.2 6.1 35.4 26.2 29.4 90.6 88.2 89.4
DeepResonance-$\beta$-Mistral (ours) 34.8 3.9 32.3 24.2 26.8 89.6 87.5 88.6
NExT-GPT-LLaMA-3 w/ Music4way 25.7 2.9 35.9 21.1 25.4 89.1 86.4 87.7
DeepResonance-$\alpha$-LLaMA-3 (ours) 34.8 5.7 37.2 25.0 29.4 90.9 88.2 89.5
DeepResonance-$\beta$-LLaMA-3 (ours) 35.4 4.5 33.9 24.8 28.0 90.1 87.8 88.9
Table 16: Results on Music4way-Any2T. The top two performances in the first block are highlighted in bold. Scores where Mistral or LLaMA-3 based models surpass their respective baselines are shown in bold. “*” denotes results that should be interpreted with caution, as the test data was included in the corresponding model’s training set.
@@
DeepResonance-$\alpha$ w/o v.c. 35.0 14.7 49.8 38.5 40.4 92.3 90.6 91.4
DeepResonance-$\alpha$ w/o m.f. 34.9 14.5 49.8 38.5 40.3 92.3 90.7 91.5
DeepResonance-$\beta$ w/o v.c. 35.2 15.1 50.4 38.9 40.7 92.4 90.7 91.5
DeepResonance-$\beta$ w/o m.f. 35.2 14.7 49.3 38.7 40.2 92.2 90.7 91.4
Table 22: Ablation study on MusicQA. The top two performances are highlighted in bold. “v.c.” and “m.f.” refer to visual captions and low-level musical features, respectively.
@@
DeepResonance-$\alpha$ w/o v.c. 25.5 2.6 23.0 21.4 21.2 87.6 86.9 87.2
DeepResonance-$\alpha$ w/o m.f. 25.9 2.7 22.9 21.6 21.3 87.6 87.0 87.3
DeepResonance-$\beta$ w/o v.c. 25.8 2.7 22.8 21.5 21.2 87.6 87.1 87.3
DeepResonance-$\beta$ w/o m.f. 25.8 2.7 22.7 21.5 21.1 87.5 87.1 87.3
Table 23: Ablation study on MusicCaps. The top two performances are highlighted in bold. “v.c.” and “m.f.” refer to visual captions and low-level musical features, respectively.
@@
DeepResonance-$\alpha$ w/o v.c. 40.7 19.7 57.2 47.7 48.1 94.0 92.6 93.2
DeepResonance-$\alpha$ w/o m.f. 40.5 19.6 56.7 48.1 48.0 93.9 92.6 93.2
DeepResonance-$\beta$ w/o v.c. 39.5 18.5 55.0 47.2 46.7 93.6 92.5 93.0
DeepResonance-$\beta$ w/o m.f. 40.0 19.1 56.3 47.7 47.5 93.8 92.5 93.1
Table 24: Ablation study on Music4way-MusicCaps. The top two performances are highlighted in bold. “v.c.” and “m.f.” refer to visual captions and low-level musical features, respectively.